OpenStudy (anonymous):

One cell phone plan costs $29.94 per month plus $.10 for MB of data. Another plan costs $32.99 per month plus $.05 for each MB of data. For how many MBs of data will the monthly bill for be the same for both plans?

OpenStudy (anonymous):

Write an equation for both. Cellphone Plan #1: y = 29.94 + 0.10x Cellphone Plan #2: y = 32.99 + 0.05x Where x = MB of data.

OpenStudy (anonymous):

Now we can graph them to see when they will be the same. Graph: https://www.desmos.com/calculator/fbs4tdtnih So what's your answer going to be? @silverwolf123

OpenStudy (anonymous):

3?